Ask your own question, for FREE!
Mathematics 6 Online
OpenStudy (anonymous):

two fair dice are rolled, what is the probability that the total showing is either even or less than four?

OpenStudy (tommynaut):

Have a go yourself first. Do you understand what "and" and "or" mean in probability questions? Usually, "and" implies multiplication (+), "or" implies addition (+). So, for example, if we have "what is the probability of rolling a 1 OR a 2 on a fair dice?", we calculate the probability of rolling a 1 and we add that to the probability of rolling a 2, because either are allowed. So, that would be 1/6 + 1/6 = 2/6 = 1/3. Try and use a similar sort of method for your problem. Work out the probability of the two rolled dice showing an even number, then add that to the probability of the dice values added together being less than 4.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
12 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
15 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!